INAF queue observing, 2020 Feb 01 - Feb 02

Observers: A. Rossi, S. Paiano

Telescope Operator: S. Allanson

Support Astronomer: B. Rothberg

Instrument: LUCI2

At sunset the sky is clear

MODS: both ok

LBC:
  • LBC Red ok; LBC blue is available although is just following the guiding of LBCred.

LUCI:
  • LUCI1: not available all week
  • LUCI2: test of AGW at beginning of the night

Temperature = 5 Humidity = 20 Wind guts >5

Summary

Night duration (12 degree - twilight) 1:30 - 13:30 UT = 12h 00m = 43200 seconds
Observing time: h min (90.3%) time given to LBTO for test of AGW: 30 min, all before 12 degree

Weather time loss min (0%) sec

Technical time loss: 70min (9.7%) sec

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Calibration flat, arcs LUCI2 zJ200 1.0" lambda=1.17   luci2.20200202.0051-0046 per Nardini
Calibration flat arcs LUCI2 K210 0.25" lambda=2.24   luci2.20200202.0047-0062 per Giannini
Gargiulo UDS_3448 / VAN003 LUCI2 zJ200 mask VAN003 1.3h luci2.20200201.0067-78 seeing >1.1 in last 4 exposures
Telluric HIP17623 LUCI2 zJ200 mask VAN003 luci2.20200201.0086-89 per Gargiulo
Rossi SN2018ibb LUCI2 JHK imaging 0.3h luci2.20200201.0099-122 completed
Giannini V900Mon LUCI2 K210 0.25" lambda=2.24 0.5/0.5h luci2.20200201.0127-156 completed
Telluric HD52266 LUCI2 K210 0.25" lambda=2.24 luci2.20200201.0161-164 per Giannini
Giannini ZCma LUCI2 K210 0.25" lambda=2.24 0.5/0.5h luci2.20200201.0181-0210 completed
Telluric HD52266 LUCI2 K210 0.25" lambda=2.24 luci2.20200201.0 again
Nardini SDSSJ0900 LUCI2 zJ200 1.0" 0.13/0.13h luci2.20200201.0223-0226 completed
Telluric HIP52849 LUCI2 zJ200 1.0" luci2.20200201.0230-0233
Nardini SDSSJ0901 LUCI2 zJ200 1.0" 0.33/0.33h luci2.20200201.02-02 completed
Annibali UGC5139 LBC g+r 2/2h completed
Nonino EGS LBC SDT_U+F972N20 2h/24 primi 2 OB
Schmidt J1149+28 LBC r+r 500s+500s    
CALIBRATION DARK LUCI2       Gargiulo, Giannini

Details

LUCI CALIBRATIONS

01:00 LBTO test on sky of AGW

01:30 LBTO test on sky finished we have the telescope! TO collimates on our field

01:38 Slew to Gargiulo VAN003; seeing 1" on DIMM
  • acq luci2.20200201.0064-66
  • OSA did the allignement because the id slits for allignement did not appeared on our RTD
  • the allignement seems good we go ahead
  • sci luci2.20200201.0067-82
  • bright star for verification is very vright inside the 3rd slit from top
  • 02:28 Seeing 1.2" on DIMM, 1" on guide camera
  • 02:45 seeing 1.1" on guider and DIMM (dimm star was bad)
  • seeing 1"-1.2" we take another hour
  • 03:10 seeing 1.2"-1.3"
  • 03:25 seeing stable above 1.3" we stop

03:27 Slew to Telluric stadard HIP 17623
  • acq luci2.20200201.0083-85
  • sci luci2.20200201.0086-89 first slit
  • acq luci2.20200201.0090-94
  • sci luci2.20200201.0095-98 second slit
  • presa 2 volet con ordine slit ribaltato

03:53 Slew to Rossi DDT imaging JHK
  • sci luci2.20200201.0099-122
  • seeing 1" on guider

04:25 Slew to Giannini V900Mon
  • acq luci2.20200201.0123-126
  • seeing 0.9"
  • sci luci2.20200201.0127-156

05:21 Slew to Telluric stadard HD52266 for Giannini
  • acq luci2.20200201.0157-160
  • sci luci2.20200201.0161-164

05:40 Slew to Giannini ZCma
  • acq luci2.20200201.0165-01
  • target no in tru slit. We create a new script as we suspect an error in the script.
  • seeing 0.7"
  • 06:15 MOS error = 54 min loss=

07:09 star agani ZCma new script
  • acq luci2.20200201.0177-0180
  • sci luci2.20200201.0181-0210
  • seeing 1.2"

07:47 Slew to Telluric stadard HD52266 for Giannini
  • acq luci2.20200201.0211-0214
  • sci luci2.20200201.0215-0218

08:05 Slew to Nardini SDSSJ0900
  • seeing 1" on guider (DIMM is not working)
  • acq luci2.20200201.0219-0222
  • sci luci2.20200201.0223-0226
  • seeing 1.1" on guider

08:28 Slew to Telluric HIP52849
  • acq luci2.20200201.0227-0229
  • 08:34 MOS error = 15 min loss =
  • sci luci2.20200201.0230-0233
  • seeing 1.0"-1.2"

08:57 Slew to Nardini SDSSJ0901
  • acq luci2.20200201.0234-0239
  • sci luci2.20200201.0238-0245
  • seeing 1" on DIMM

09:35 Seeing 0.8-1" switch to LBC

09:57 Slew to Annibali R14
  • DIMM seeing 0.9"
  • dohybrid
  • copointing
  • 10:18 send UGC5139_1
  • seeing 0.8" on image
  • collimation still good after 20 min
  • 10:37 send UGC5139_2
  • 10:45 DIMM 1.0" and on image
  • dofpia
  • 11:06 send UGC5139_3
  • seeing 0.9" DIMM and image

11:32 Slew to Nonino EGS
  • DIMM seeing 0.9"
  • dofpia
  • copointing
  • 11:44 send IT-2019B-017_1.xml
  • FWHM 0.7" on LBCr at 11:50
  • FWHM 1.0" on LBCb at 11:50
  • dofpia
  • 12:26 send IT-2019B-017_2.xml
  • DIMM seeing 0.9"
  • DIMM seeing 0.8" at 12:50

13:02 Slew to Schmidt J1149+28
  • dofpia
  • copointing
  • send OB missing r run 60: blue+red_r+r.xml
  • DIMM seeing 0.8"
  • done only 5+5 images, then counts too high

LBC SKYFLATS SDT_U+F972N20

LUCI DARKS
  • Gargiulo 300.0x1 MER
  • Gargiulo 3.0x5 INT
  • Giannini darks 5x12.0, 6x5.0 LIR , 2x30.0 LIR 10x2.5 LIR
Topic revision: r9 - 07 Feb 2020, AndreaRossi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback